# AGN - host galaxy photometric decomposition using a fast, accurate and precise deep learning approach

### v2

@@ 3.1 Mock AGN injection
For each galaxy, five different images were created with five different AGN contribution fractions, chosen randomly from the 20 possible discrete AGN contribution fractions. Example images of these simulated galaxies (from bright to faint) with varying levels of AGN contribution can be seen in Fig. 2. Visually, it is clear that it can be very difficult to discern the presence of the host galaxy when the AGN contribution fraction is large, particularly for fainter and higher-redshift objects. Later we will examine how well we can recover $f_{AGN}$ as a function of the intrinsic AGN contribution fraction, redshift, galaxy size, S/N, etc, using a set of common performance metrics.
@@ 3.2 Deep learning CNN
For this work, we chose a ConvNeXt-Base architecture, which was pre-trained on the Galaxy Zoo dataset of over 820k images and 100 million volunteer votes to morphological questions. In order to perform a regression task (as is the case in this study), we added a linear head with a sigmoid function (to restrict the output to be between 0 and 1) and used a mean square error loss function to train the network. A diagram of the ConvNeXt-Base network can be seen in Fig. 3. For this work, we retrained the last two blocks of the network and the linear head, while the rest of the network parameters were frozen to keep the optimal values found for the pre-trained data from Galaxy Zoo.
@@ 3.3 GALFIT
Figure 5: Sérsic + PSF decomposition of four example galaxies (with AGN contributon fraction varying from high to low from top to bottom) on which we performed GALFIT. Images of the original galaxy, the model (Sérsic + PSF), and the residual (original--model) are shown from left to right. Images are 3.84″across, and are displayed with an inverse arcsinh scaling.
@@
$\sum(R)=\sum_{e}\left\{-\kappa_{n}\left[\left(\frac{R}{R_{e}}\right)^{1/n}-1\right]\right\},$ (9)
@@
We ran GALFIT with a combination of a single Sérsic profile and a PSF model for the main central galaxy in each image. Neighbouring galaxies were masked, unless their light overlaps with the central galaxy, in which case they were fitted simultaneously with a Sérsic profile. The initial estimates of the model parameters can have an impact on whether GALFIT finds a good fit or not. That is why we ran GALFIT with different initial parameters of Sérsic index and magnitudes. We chose the best model to be the one with the lowest reduced $\chi^{2}$. For the Sérsic index of the main galaxy, we chose as initial values $n=1,2,4$, alternatively. For the magnitudes of the Sérsic and PSF models, we used three different combinations of the initial parameters. In the first scenario, both models were set to be equal to a magnitude that corresponds to half of the total flux obtained from Sextractor. In the other two scenarios, the magnitude of the PSF (Sérsic) model was set to be 80% of the total flux while the magnitude of the Sérsic (PSF) model was set to be 20%. For the rest of the parameters (position x and y, axis ratio, position angle and effective radius), we used as initial values those obtained by Sextractor. For the neighbouring galaxies that were fitted simultaneously, we also used the Sextractor parameters plus a Sérsic index of $n=2$.
@@ 4.1 Zoobot model performance
Figure 6: Comparison between the real injected AGN contribution fraction and the AGN contribution fraction obtained from the Zoobot model on the subset of 4800 galaxies across the whole redshift range ($30.5<z<$ 3). The comparison shows a mean difference between the two quantities ($\Delta f_{AGN}=f_{AGN}$ [Injected] $-f_{AGN}$ [Zoobot]) of $-0.0018$ and an overall RMSE = 0.013. The solid diagonal line is the 1:1 line. The top plot shows the mean difference and its dispersion as a function of the injected AGN contribution fraction. The colour bar indicates the number of sources in each bin.
Table 1: Performance statistics for Zoobot and GALFIT.
@@
Zoobot $0.013\pm 0.005$ $0.076\pm 0.005$ $-0.0018\pm 0.0002$ $0.065\pm 0.004$ $0.061\pm 0.003$
ZoobotM $0.014\pm 0.006$ $0.102\pm 0.009$ $-0.0017\pm 0.0005$ $0.072\pm 0.010$ $0.065\pm 0.010$
ZoobotNM $0.013\pm 0.006$ $0.073\pm 0.005$ $-0.0018\pm 0.0002$ $0.065\pm 0.004$ $0.061\pm 0.004$
GALFIT ($f_{host}$ from original image) $0.12\pm 0.07$ $0.19\pm 0.01$ $-0.018\pm 0.002$ $0.19\pm 0.01$ $0.13\pm 0.01$
GALFITM ($f_{host}$ from original image) $0.13\pm 0.07$ $0.27\pm 0.05$ $-0.024\pm 0.004$ $0.25\pm 0.03$ $0.18\pm 0.02$
GALFITNM ($f_{host}$ from original image) $0.12\pm 0.07$ $0.17\pm 0.01$ $-0.016\pm 0.002$ $0.16\pm 0.01$ $0.11\pm 0.01$
GALFIT ($f_{host}$ from model) $0.17\pm 0.08$ $0.51\pm 0.02$ $-0.075\pm 0.003$ $0.46\pm 0.01$ $0.35\pm 0.01$
GALFITM ($f_{host}$ from model) $0.16\pm 0.09$ $0.48\pm 0.07$ $-0.087\pm 0.007$ $0.54\pm 0.04$ $0.39\pm 0.04$
GALFITNM ($f_{host}$ from model) $0.17\pm 0.08$ $0.52\pm 0.02$ $-0.073\pm 0.003$ $0.45\pm 0.02$ $0.34\pm 0.01$
 [^3] We summarise the overall performance from Zoobot and GALFIT (for both ways of calculating the AGN fraction contribution), in terms of the RMSE, RAE, mean difference and outlier fractions (at different percentage levels, 20% and 30%). We show the results for the whole sample and mergers (M) and non-mergers (NM) separately.
@@
which measures the average difference between the predicted values from a model (e.g. Zoobot or GALFIT) and the actual injected values. In Fig. 6 we show the predicted values of $f_{AGN}$, obtained from Zoobot, versus the real $f_{AGN}$ for all redshifts, which shows a very tight correlation across the whole dynamic range. Quantitatively, we found an overall value of $\text{RMSE}=0.013$ for the Zoobot model, demonstrating very good recovery (both in terms of accuracy and precision) of the injected real AGN contribution fractions. In addition, the top panel of Fig. 6 shows the mean and the dispersion of the difference between the real and predicted values ($\Delta f_{AGN}=f_{AGN}$ [injected] $-f_{AGN}$ [Zoobot]) as a function of the injected AGN fraction, with a value for the mean of $\langle{\Delta f_{AGN}}\rangle=-0.0018$, and dispersion of $\sigma(\Delta f_{AGN})=0.013$. This shows that there is very little bias from the Zoobot predictions to the actual values, and it is independent of $f_{AGN}$, as the difference between the real and predicted value over the whole range of $f_{AGN}$ is always below -0.0043. To check if the performance depends on galaxy structural properties, we further explore whether the Sérsic index impacts how well Zoobot can predict the AGN contribution fraction (see Fig. 15 in Appendix A) and find that for galaxies with $n<1$ the RMSE increases by 30% (compared to the overal RMSE of 0.013), to RMSE$=0.017$ and that high Sérsic indexes ($n>6$) do not lead to worse predictions (i.e., similar RMSE to the overal RMSE of $0.013$).
@@
Figure 9: Outlier fraction (calculated as the fraction of galaxies with RAE $>20\%$) as a function of the injected AGN contribution fraction (top left panel), redshift (top right), S/N (bottom left) and $r$ (bottom right). The purple lines correspond to the results from Zoobot and the green, from GALFIT. The solid lines correspond to the whole sample while the dashed and dotted lines correspond to the merger and non-merger galaxies, respectively.
In Fig. 7 we show the RMSE of the Zoobot predictions (purple lines) as a function of the injected AGN contribution fraction, redshift, S/N and kron radius ($r_{kron}$, calculated with SExtractor in the F150W filter). The solid lines show the RMSE of the whole test subset. To separately investigate how the performance might change for highly disturbed or irregular galaxies, we also show galaxies classified as mergers (dashed lines) and galaxies classified as non-mergers (dotted lines), separately. In general, the RMSE is slightly lower for non-merging galaxies, but the difference with mergers is very small. The RMSE is larger for lower values of the injected AGN contribution fraction and reaches a maximum mean value of 0.020. At $f_{AGN}>0.2$, Zoobot achieves smaller errors than the coefficient of variation of the PSF, that is, smaller error than the error on the PSF resulting from the intrinsic temporal and spatial variations (shaded region of Fig. 7). In other words, the precision of our DL-based method is better than the intrinsic variations of the observed PSF. This is only possible because the DL method is trained on data which include the full range of observed PSFs. There is an increase in the RMSE with increasing redshift, with a maximum value of RMSE $=0.026$ at $z=3$, possibly because galaxies at higher redshifts are smaller and tend to have lower S/N (i.e. more dominated by noise). Indeed, we see that generally the RMSE increases with decreasing $r_{kron}$ and S/N.
In Fig. 8 we show the relative absolute error (RAE), which is the ratio between the absolute error divided by the real value [^4] The relative error is not well defined when the real AGN fraction is equal to zero. In order to calculate the relative error when $f_{AGN}$ [injected] $=0$, we approximate it with $f_{AGN}$ [injected] $=1e-3$.,
@@ 4.3 Application to JWST galaxies
Finally, we apply the trained DL model to our stellar mass complete sample of real JWST galaxies as described in Section 2.1. In Fig. 12, we show the distribution of the DL predicted AGN contribution fractions. Considering the overal RMSE of our DL-based method in recovering the intrinsic $f_{AGN}$ is 0.013, a $5\sigma$ cut corresponding to $f_{AGN}=0.065$ can be used to select galaxies for which the method is confident in identifying a non-negligible contribution from a central point source component. Indeed, the distribution shown in Fig.12 clearly separates at around this $5\sigma$ cut value. Note that many traditional methods of selecting AGN use a binary approach to separate galaxies into those which host AGN and those which do not. In reality, there is a continuous distribution of AGN contribution fraction. Therefore, although a binary AGN vs non-AGN selection approach is useful in identifying galaxies with dominant AGN, it is limited in the sense that it does not reflect the full spectrum. Adopting the $5\sigma$ cut at $f_{AGN}=0.065$, we find a total of $3337\begin{subarray}{c}+218\\ -196\end{subarray}$ galaxies hosting a non-negligile level of AGN activity in the 0.28 deg2 of the COSMOS field analysed in this paper. Therefore, with respect to the total size of our stellar mass complete sample, 13% of the galaxies are identified by our method to have a non-neglible level of AGN contribution in the JWST/NIRCAM F150W filter.
@@
Using the continuous parameter $f_{AGN}$ predicted by our DL method, users can choose different cuts on AGN contribution fraction to construct an AGN sample depending on the specific requirements. For example, we can classify galaxies as AGN based on the fractional AGN contribution, using a cut at $f_{AGN}>0.2$, which leads to $937\begin{subarray}{c}+111\\ -170\end{subarray}$ galaxies being classified as AGN over 0.28 deg2. With a less conservative cut at $f_{AGN}>0.1$, we find $2611\begin{subarray}{c}+358\\ -753\end{subarray}$ AGN within our sample. Furthermore, we find that there are 18 galaxies with $f_{AGN}>0.7$ (shown in Fig. 17 in the appendix). There are five galaxies with spectroscopic information within those 18 galaxies. Four out of those five (i.e. 80%) show optical broad lines in their spectra, as indicated by the specific flag in the release catalogue of spectroscopic redshifts available to the COSMOS collaboration from KECK follow up observations (Khostovan et al 2025, in prep). Based on this, we conclude that most galaxies with high $f_{AGN}$, as predicted by our method, are consistent with optical spectroscopically confirmed AGN (when that information is available)
@@
Next, we compare our AGN based on the estimated $f_{\rm AGN}$ with AGN samples selected via X-ray detection or MIR colours. Using the cut at $f_{\rm AGN}>0.2$, we find that 20 % of the X-ray AGN are also selected as AGN based on the estimated $f_{\rm AGN}$, while 8 % of the MIR selected AGN are also selected by our method. If we use a less conservative cut at $f_{\rm AGN}>0.1$, then the overlapping fractions increase for both AGN selections, to 15 % overlap with the MIR-selected AGN and 33 % overlap with the X-ray AGN. There is, however, a large fraction of the X-ray and MIR AGN with very low AGN contribution fraction $f_{\rm AGN}<0.1$, and therefore, are not selected as AGN by our DL method based on a 0.1 threshold on $f_{AGN}$. This is to be expected, as it is well-known that some X-ray and MIR AGN are not detected by optical diagnostics. This phenomena could be explained by a bright host galaxy that dilutes the light of the AGN or by the presence of copious dust in the host galaxy, particularly when the galaxy is viewed edge-on, causing heavy obscuration (Satyapal et al., 2008; Jackson et al., 2012; Smith et al., 2014; Hickox & Alexander, 2018; Fitriana & Murayama, 2022). In Fig. 18 and 19 we show a random subset of the X-ray and MIR AGN with $f_{\rm AGN}<0.065$ (i.e., below the $5\sigma$ cut based on the overall RMSE of our DL method), respectively. Many of these galaxies display spiral, clumpy or edge-on morphologies, in which higher dust content is expected. And in all cases, they do not show any obvious nuclear component in the JWST/NIRCAM F150W images, which explains the low predicted $f_{\rm AGN}$ despite being classified as X-ray or MIR AGN.
@@
Figure 14: Top panel: Overlapping fraction of AGN also identified by our DL method as a function of the adopted cut on the X-ray or the rest-frame 6 $\mu$m luminosity for the X-ray and MIR-selected AGN. The solid lines show the overlapping fraction by applying a cut at $f_{\rm AGN}>0.2$, while the dashed lines show the overlapping fraction if we select AGN by requiring $f_{\rm AGN}>0.1$. Bottom panel: AGN luminosity in the JWST/NIRCAM F150W filter versus the X-ray luminosity or the rest-frame 6 $\mu m$ luminosity, for the X-ray-selected or MIR-selected AGN. The dash-dotted lines show the respective linear fits.

### v3

@@ 3.1 Mock AGN injection
For each galaxy, five different images were created with five different AGN contribution fractions, chosen randomly from the 20 possible discrete AGN contribution fractions. Example images of these simulated galaxies (from bright to faint) with varying levels of AGN contribution can be seen in Fig. 2. Visually, it is clear that it can be very difficult to discern the presence of the host galaxy when the AGN contribution fraction is large, particularly for fainter and higher-redshift objects. Later, we will examine how well we can recover $f_{AGN}$ as a function of the intrinsic AGN contribution fraction, redshift, galaxy size, S/N, etc, using a set of common performance metrics.
@@ 3.2 Deep learning CNN
For this work, we chose a ConvNeXt-Base architecture, which was pre-trained on the Galaxy Zoo dataset of over 820k images and 100 million volunteer votes on morphological questions. In order to perform a regression task (as is the case in this study), we added a linear head with a sigmoid function (to restrict the output to be between 0 and 1) and used a mean square error loss function to train the network. A diagram of the ConvNeXt-Base network can be seen in Fig. 3. For this work, we retrained the last two blocks of the network and the linear head, while the rest of the network parameters were frozen to keep the optimal values found for the pre-trained data from Galaxy Zoo. Our model was trained on a v100 GPU and took 72 hours to complete. Once the model is trained, it takes Zoobot $6\times 10^{-3}$ seconds to predict one galaxy.
@@ 3.3 GALFIT
Figure 5: Sérsic + PSF decomposition of four example galaxies (with AGN contributon fraction varying from high to low from top to bottom) on which we performed GALFIT. Images of the original galaxy, the model (Sérsic + PSF), and the residual (original-model) are shown from left to right. Images are 3.84″across, and are displayed with an inverse arcsinh scaling.
@@
$\sum(R)={\sum}_{e}exp\left\{-\kappa_{n}\left[\left(\frac{R}{R_{e}}\right)^{1/n}-1\right]\right\},$ (10)
@@
We ran GALFIT with a combination of a single Sérsic profile and a PSF model for the main central galaxy in each image. Neighbouring galaxies were masked, unless their light overlaps with the central galaxy, in which case they were fitted simultaneously with a Sérsic profile. The initial estimates of the model parameters can have an impact on whether GALFIT finds a good fit or not. That is why we ran GALFIT with different initial parameters of Sérsic index and magnitudes. We chose the best model to be the one with the lowest reduced $\chi^{2}$. For the Sérsic index of the main galaxy, we chose as initial values $n=1,2,4$, alternatively. For the magnitudes of the Sérsic and PSF models, we used three different combinations of the initial parameters. In the first scenario, both models were set to be equal to a magnitude that corresponds to half of the total flux obtained from Sextractor. In the other two scenarios, the magnitude of the PSF (Sérsic) model was set to be 80% of the total flux while the magnitude of the Sérsic (PSF) model was set to be 20%. For the rest of the parameters (position x and y, axis ratio, position angle and effective radius), we used as initial values those obtained by Sextractor. In total, this resulted in 9 different model configurations (3 Sérsic index values $\times$ 3 magnitude combinations), from which we selected the one with the lowest reduced $\chi^{2}$ as the final best fit. For the neighbouring galaxies that were fitted simultaneously, we also used the Sextractor parameters plus a Sérsic index of $n=2$. On average, GALFIT took 15 seconds per galaxy to complete the fitting procedure (2500 times slower than the prediction time from Zoobot), which resulted in a total of 90 hours to fill all 9 model configurations (resulting from the different combinations of initial Sérsic indices and magnitudes) to the 2400 galaxies in our test subset.
@@ 4.1 Zoobot model performance
Figure 6: Comparison between the real injected AGN contribution fraction and the AGN contribution fraction obtained from the Zoobot model on the subset of 4800 galaxies across the whole redshift range ($0.5<z<3$). The comparison shows a mean difference between the two quantities ($\Delta f_{AGN}=f_{AGN}$ [Injected] $f_{AGN}$ [Zoobot]) of 0.0018-0.0018 and an overall RMSE = 0.013. The solid diagonal line is the 1:1 line. The top plot shows the mean difference and its dispersion as a function of the injected AGN contribution fraction. The colour bar indicates the number of sources in each bin.
Table 1: Performance statistics for Zoobot and GALFIT. [^3] We summarise the overall performance from Zoobot and GALFIT (for both ways of calculating the AGN fraction contribution), in terms of the RMSE, RAE, mean difference and outlier fractions (at different percentage levels, 20% and 30%). We show the results for the whole sample and mergers (M) and non-mergers (NM) separately.
@@
Zoobot $013\pm 0.005 0.0$$76\pm 0.005 0.0$$018\pm 0.0002 0.0$$65\pm 0.004 0.0$$061\pm 0.003$
ZoobotM $0.014\pm 0.006$ $102\pm 0.009 0.0$$017\pm 0.0005 0.0$$72\pm 0.010 0.0$$065\pm 0.010$
ZoobotNM $013\pm 0.006 0.0$$73\pm 0.005 0.0$$018\pm 0.0002 0.0$$65\pm 0.004 0.0$$061\pm 0.004$
GALFIT ($f_{host}$ from original image) $0.12\pm 0.07$ $19\pm 0.01 0.0$$018\pm 0.002$ $0.19\pm 0.01$ $0.13\pm 0.01$
GALFITM ($f_{host}$ from original image) $0.13\pm 0.07$ $27\pm 0.05 0.0$$024\pm 0.004$ $0.25\pm 0.03$ $0.18\pm 0.02$
GALFITNM ($f_{host}$ from original image) $0.12\pm 0.07$ $17\pm 0.01 0.0$$016\pm 0.002$ $6\pm 0.01 0.11±0$$1\pm 0.01$
GALFIT ($f_{host}$ from model) $0.17\pm 0.08$ $51\pm 0.02 0.0$$075\pm 0.003$ $0.46\pm 0.01$ $0.35\pm 0.01$
GALFITM ($f_{host}$ from model) $0.16\pm 0.09$ $48\pm 0.07 0.0$$087\pm 0.007$ $0.54\pm 0.04$ $0.39\pm 0.04$
GALFITNM ($f_{host}$ from model) $0.17\pm 0.08$ $52\pm 0.02 0.0$$073\pm 0.003$ $0.45\pm 0.02$ $0.34\pm 0.01$
@@
which measures the average difference between the predicted values from a model (e.g. Zoobot or GALFIT) and the actual injected values. In Fig. 6 we show the predicted values of $f_{AGN}$, obtained from Zoobot, versus the real $f_{AGN}$ for all redshifts, which shows a very tight correlation across the whole dynamic range. Quantitatively, we found an overall value of $\text{RMSE}=0.013$ for the Zoobot model, demonstrating very good recovery (both in terms of accuracy and precision) of the injected real AGN contribution fractions. In addition, the top panel of Fig. 6 shows the mean and the dispersion of the difference between the real and predicted values ($\Delta f_{AGN}=f_{AGN}$ [injected] $f_{AGN}$ [Zoobot]) as a function of the injected AGN fraction, with a value for the mean of $\langle{\Delta f_{AGN}}\rangle=-0.0018$, and dispersion of $\sigma(\Delta f_{AGN})=0.013$. This shows that there is very little bias from the Zoobot predictions to the actual values, and it is independent of $f_{AGN}$, as the difference between the real and predicted value over the whole range of $f_{AGN}$ is always below -0.0043. To check if the performance depends on galaxy structural properties, we further explore whether the Sérsic index impacts how well Zoobot can predict the AGN contribution fraction (see Fig. 15 in Appendix A) and find that for galaxies with $n<1$ the RMSE increases by 30% (compared to the overal RMSE of 0.013), to RMSE=0.017=0.017 and that high Sérsic indexes ($n>6$) do not lead to worse predictions (i.e., similar RMSE to the overal RMSE of $0.013$).
@@
Figure 9: Outlier fraction (calculated as the fraction of galaxies with RAE $20\%$) as a function of the injected AGN contribution fraction (top left panel), redshift (top right), S/N (bottom left) and $r$ (bottom right). The purple lines correspond to the results from Zoobot and the green, from GALFIT. The solid lines correspond to the whole sample while the dashed and dotted lines correspond to the merger and non-merger galaxies, respectively.
In Fig. 7 we show the RMSE of the Zoobot predictions (purple lines) as a function of the injected AGN contribution fraction, redshift, S/N and kron radius ($r_{kron}$, calculated with SExtractor in the F150W filter). The solid lines show the RMSE of the whole test subset. To separately investigate how the performance might change for highly disturbed or irregular galaxies, we also show galaxies classified as mergers (dashed lines) and galaxies classified as non-mergers (dotted lines), separately. In general, the RMSE is slightly lower for non-merging galaxies, but the difference with mergers is very small. The RMSE is larger for lower values of the injected AGN contribution fraction and reaches a maximum mean value of 0.020. At $f_{AGN}>0.2$, Zoobot achieves smaller errors than the coefficient of variation of the PSF, that is, smaller error than the error on the PSF resulting from the intrinsic temporal and spatial variations (shaded region of Fig. 7). In other words, the precision of our DL-based method is better than the intrinsic variations of the observed PSF. This is only possible because the DL method is trained on data which include the full range of observed PSFs. There is an increase in the RMSE with increasing redshift, with a maximum value of RMSE =0.026=0.026 at $z=3$, possibly because galaxies at higher redshifts are smaller and tend to have lower S/N (i.e. more dominated by noise). Indeed, we see that generally the RMSE increases with decreasing $r_{kron}$ and S/N. However, we also observe a small rise in RMSE at the highest S/N end, which is likely due to low number statistics in that bin; in such cases, even a single outlier can have a disproportionately large effect on the RMSE.
In Fig. 8 we show the relative absolute error (RAE), which is the ratio between the absolute error divided by the real value [^4] The relative error is not well defined when the real AGN fraction is equal to zero. In order to calculate the relative error when $f_{AGN}$ [injected] =0=0, we approximate it with $f_{AGN}$ [injected] =1e3=1e-3.,
@@ 4.3 Application to JWST galaxies
Finally, we apply the trained DL model to our stellar mass complete sample of real JWST galaxies as described in Section 2.1. In Fig. 12, we show the distribution of the DL predicted AGN contribution fractions. Considering the overal RMSE of our DL-based method in recovering the intrinsic $f_{AGN}$ is 0.013, a $5\sigma$ cut corresponding to $f_{AGN}=0.065$ can be used to select galaxies for which the method is confident in identifying a non-negligible contribution from a central point source component. Indeed, the distribution shown in Fig.12 clearly separates at around this $5\sigma$ cut value.
It is important to note, however, that the AGN contribution fraction measures the nuclear light excess and may not necessarily be interpreted as exclusively tracing AGN emission, since compact star-forming regions, stellar clusters, or other unresolved components, particularly at high redshift, could also partially contribute to the detected central point-source signal. Additional multi-wavelength and/or spectroscopic diagnostics are therefore needed to confirm or refute the AGN nature of these candidates.
Many traditional methods of selecting AGN use a binary approach to separate galaxies into those which host AGN and those which do not. In reality, there is a continuous distribution of AGN contribution fraction. Therefore, although a binary AGN vs non-AGN selection approach is useful in identifying galaxies with dominant AGN, it is limited in the sense that it does not reflect the full spectrum. Adopting the $5\sigma$ cut at $f_{AGN}=0.065$, we find a total of $3337\begin{subarray}{c}+218\\ -196\end{subarray}$ galaxies hosting a non-negligile level of AGN activity in the 0.28 deg2 of the COSMOS field analysed in this paper. The uncertainties correspond to varying the threshold by the RMSE of our method (0.013), i.e., counting galaxies with $f_{AGN}>0.065\pm 0.013$, which yields lower and upper thresholds of 0.052 and 0.078 respectively. Therefore, with respect to the total size of our stellar mass complete sample, 13% of the galaxies are identified by our method to have a non-neglible level of AGN contribution in the JWST/NIRCam F150W filter.
@@
Using the continuous parameter $f_{AGN}$ predicted by our DL method, users can choose different cuts on AGN contribution fraction to construct an AGN sample depending on the specific requirements. For example, we can classify galaxies as AGN based on the fractional AGN contribution, using a cut at $f_{AGN}>0.2$, which leads to $937\begin{subarray}{c}+111\\ -170\end{subarray}$ galaxies being classified as AGN over 0.28 deg2. With a less conservative cut at $f_{AGN}>0.1$, we find $2611\begin{subarray}{c}+358\\ -753\end{subarray}$ AGN within our sample. Furthermore, we find that there are 18 galaxies with $f_{AGN}>0.7$ (shown in Fig. 17 in the appendix). There are five galaxies with spectroscopic information within those 18 galaxies. Four out of those five (i.e. 80%) show optical broad lines in their spectra, as indicated by the specific flag in the release catalogue of spectroscopic redshifts available to the COSMOS collaboration from KECK follow up observations (Khostovan et al 2025, in prep). Based on this, we conclude that most galaxies with high $f_{AGN}$, as predicted by our method, are consistent with optical spectroscopically confirmed AGN (when that information is available).
To further investigate the nature of our $f_{AGN}$-selected galaxies, we perform a complementary analysis presented in Appendix C. Because the JWST/NIRCam F150W filter probes different rest-frame wavelengths across our redshift range, we train additional networks using the other three NIRCam filters available in COSMOS-Web (F115W, F277W and F [^4] W) to partially mitigate redshift effect and provide additional colour information. In Appendix C.3, we examine the colours of AGN and non-AGN galaxies, as well as MIR-selected AGN. We find that our $f_{AGN}$-selected sources display systematically redder colours than non-AGN, in the same direction as the traditionally selected MIR AGN, which may be consistent with emission from dust heated by AGN.
@@
Next, we compare our AGN based on the estimated $f_{\rm AGN}$ with AGN samples selected via X-ray detection or MIR colours. Using the cut at $f_{\rm AGN}>0.2$, we find that 20 % of the X-ray AGN are also selected as AGN based on the estimated $f_{\rm AGN}$, while 8 % of the MIR selected AGN are also selected by our method. If we use a less conservative cut at $f_{\rm AGN}>0.1$, then the overlapping fractions increase for both AGN selections, to 15 % overlap with the MIR-selected AGN and 33 % overlap with the X-ray AGN. There is, however, a large fraction of the X-ray and MIR AGN with very low AGN contribution fraction $f_{\rm AGN}<0.1$, and therefore, are not selected as AGN by our DL method based on a 0.1 threshold on $f_{AGN}$. This is to be expected, as it is well-known that some X-ray and MIR AGN are not detected by optical diagnostics (Satyapal et al., 2008; Smith et al., 2014; LaMassa et al., 2019; Comerford et al., 2022). This phenomenon could be explained by a bright host galaxy that dilutes the light of the AGN or by the presence of copious dust in the host galaxy, particularly when the galaxy is viewed edge-on, causing heavy obscuration (Satyapal et al., 2008; Jackson et al., 2012; Smith et al., 2014; Hickox & Alexander, 2018; Fitriana & Murayama, 2022). In Fig. 18 and 19 we show a random subset of the X-ray and MIR AGN with $f_{\rm AGN}<0.065$ (i.e., below the $5\sigma$ cut based on the overall RMSE of our DL method), respectively. Many of these galaxies display spiral, clumpy or edge-on morphologies, in which higher dust content is expected. And in all cases, they do not show any obvious nuclear component in the JWST/NIRCAM F150W images, which explains the low predicted $f_{\rm AGN}$ despite being classified as X-ray or MIR AGN. A similar analysis using the same methodology was done on Euclid/VIS imaging (Euclid Collaboration: Margalef-Bentabol et al., 2025), where the overlaps with X-ray and MIR AGN selections were found to be higher. However, such variations are expected, as the resulting overlap fractions depend sensitively on several factors, including the intrinsic luminosity of the AGN population, the rest-frame wavelength range probed by each dataset, and the specific selection criteria.
We also computed the fraction of our AGN, based on $f_{\rm AGN}$ that are found in the X-ray and MIR-selected AGN samples. For sources with $f_{\rm AGN}>0.2$, we find that 6 % are also X-ray AGN and 5 % are MIR AGN. These fractions decrease to 4 % and 3 %, respectively, when considering a lower threshold of $f_{\rm AGN}>0.1$. This trend suggests that X-ray and MIR selection methods preferentially identify more powerful AGN, while our optical method is sensitive to a broader population, including less dominant or possibly obscured AGN. This result reinforces the notion that different AGN selection techniques probe different AGN populations. We also note that an $f_{\rm AGN}$ value above the selection threshold does not necessarily indicate true AGN activity; in some cases, the $f_{\rm AGN}$ signal may arise from central sources unrelated to AGN. We also investigate whether the $f_{\rm AGN}$ distributions differ for galaxies identified as X-ray or MIR AGN. As shown in Fig. 12, the distributions for these subsamples closely follow that of the whole sample, indicating that our method predicts similar $f_{\rm AGN}$ values regardless of whether the galaxy is independently classified as an AGN in X-ray or MIR. This suggests that the low overlap is not due to a bias in how our method treats X-ray or MIR AGN, but rather reflects the differences in the AGN populations each method is sensitive to.
@@
[…]